OpenStudy (anonymous):

Find the inverse of f(x) = x + 5.

OpenStudy (anonymous):

y=x+5 x=y+5 x-5=y

OpenStudy (anonymous):

you switch x and y and then solve for y

OpenStudy (anonymous):

So what's the answer?

OpenStudy (anonymous):

f^-1(x)=x-5 the notation for the inverse of f(x) is f^-1(x)

OpenStudy (anonymous):

The inverse of y = 5 − 3x is y equals negative one-third x plus five-thirds?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The inverse of y = 5 − 3x is y equals negative one-third x plus five-thirds?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The inverse of f of x equals one half x plus 2 is f-1(x) = 2x − 4?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

This ?

OpenStudy (anonymous):

A y>=3

OpenStudy (anonymous):

This one and thank you so much

OpenStudy (anonymous):

B you're welcome

OpenStudy (anonymous):

And these questions

OpenStudy (anonymous):

False, B, D

OpenStudy (anonymous):

I still need more help

OpenStudy (anonymous):

I would but i have to go someone is helping me with my problem. i can help you later though

OpenStudy (anonymous):

Okay how tho?